# Torsion of $\mathbb{Q}$-curves over number fields of small odd prime degree

### v1

@@ 3. Torsion over cubic number fields
[…]
@@
To prove Theorem 1.8, it remains to prove that none of the following six groups appears as torsion of a $\mathbb{Q}$-curve over a cubic field.
[…]
@@ 3.1. Eliminating $\mathbb{Z}/11\mathbb{Z}$
[…]
@@
There are two rational $j$-invariants for which the corresponding elliptic curve $E/\mathbb{Q}$ has an $11$-isogeny over $\mathbb{Q}$. For a fixed $j$-invariant, any two rational elliptic curves with this $j$-invariant are quadratic twists. This means that we can pick any rational elliptic curve with this $j$-invariant and check that its Galois image doesn’t have the properties it would need to have by Corollary 2.4.
The $j$-invariants for which there is an $11$-isogeny over $\mathbb{Q}$ are $-11^{2}$ and $-11\cdot 131^{3}$. Examples of elliptic curves with these $j$-invariant are 1089.c1 and 1089.c2 from LMFDB.
For $E^{\prime\prime}$ being any of these curves, the image of the mod 11 Galois representation $\rho_{E^{\prime\prime},11}(G_{\mathbb{Q}})$ has $220$ elements. Since $3\nmid 220$, we have $\rho_{E^{\prime\prime},11}(G_{K})=\rho_{E^{\prime\prime},11}(G_{\mathbb{Q}})$.
Inspecting the images $\rho_{E^{\prime\prime},11}(G_{\mathbb{Q}})$, it turns out that one can find matrices which have neither $1$ nor $-1$ as an eigenvalue, so we reach a contradiction with Corollary 2.4.
$\Box$
@@ 3.2. Eliminating $\mathbb{Z}/15\mathbb{Z}$
[…]
@@
We will look at curves from the isogeny class 1600.i on LMFDB. Note that $-I$ is contained in the mod $15$ image.
By Lemma 2.1, we have $\rho_{E^{\prime},15}\sim\psi\cdot\rho_{E^{\prime\prime},15}$, where $\psi$ is some quadratic character. Furthermore, any element of $\rho_{E^{\prime},15}(G_{K})$ has $1$ as an eigenvalue when reduced modulo $3$ and modulo $5$. This means that any element of $\rho_{E^{\prime\prime},15}(G_{K})$ either has $1$ as an eigenvalue modulo $3$ and modulo $5$, or it has $-1$ as an eigenvalue modulo $3$ and modulo $5$. Now, for each of the four curves $E^{\prime\prime}$ from the class 1600.i, we do the following:
@@
  • In each of the subgroups $H$ from the list, find a matrix $A$ such that at least one matrix among $A\pmod{3}$, $A\pmod{5}$ doesn’t have $1$ as an eigenvalue, and at least one matrix among $A\pmod{3}$, $A\pmod{5}$ doesn’t have $-1$ as an eigenvalue.
@@
      The matrices $\begin{bmatrix}12&1\\ 10&6\end{bmatrix}$ and $\begin{bmatrix}6&5\\ 5&6\end{bmatrix}$ are in $\rho_{E^{\prime\prime},15}(G_{K})$. They don’t have $-1$ as eigenvalues mod $5$, so they themselves have to be in $\rho_{E^{\prime},15}(G_{K})$. However, their product equals $-I$ mod $3$, which doesn’t have $1$ as an eigenvalue mod $3$, so we reach a contradiction.
@@
      Similarly as in the previous case, we end up with $-I$ in $\rho_{E^{\prime},15}(G_{K})$, which is a contradiction.
@@ 3.3. Eliminating $\mathbb{Z}/16\mathbb{Z}$
Proof.   Suppose that some $\mathbb{Q}$-curve $E$ over a cubic field $K$ has a point $P$ of order $16$. We have the following situation:
[…]
@@
Here $E^{\prime}$ is an elliptic curve over $K$ with rational $j$-invariant, and $E^{\prime\prime}$ is a quadratic twist of $E^{\prime}$ which is defined over $\mathbb{Q}$. We use Corollary 2.6. The mod $16$ image $\rho_{E^{\prime},16}(G_{K})$ is conjugate to a subgroup of the group of all matrices of the form
$\begin{bmatrix}x&2^{4-s}u\\ 2^{s}v&1\end{bmatrix},$
for some fixed $s$ satisfying $0\leq s\leq 4$.
Note that $E^{\prime\prime}$ has at least a $4$-isogeny over a cubic extension.
If $s\neq 2$, it even has an $8$-isogeny. It follows from [17, Proposition 3.2.] that the $8$-isogeny is defined over $\mathbb{Q}$. However, the mod $16$ image $\rho_{E^{\prime\prime},16}(G_{\mathbb{Q}})$ is then conjugate to a subgroup of the Borel subgroup when viewed modulo $2$, so its order is not divisible by $3$. Thus, $\rho_{E^{\prime\prime},16}(G_{\mathbb{Q}})=\rho_{E^{\prime\prime},16}(G_{K})$.
@@
Now suppose that $s=2$. Then $E^{\prime\prime}$ has a $4$-isogeny over a cubic extension. If this $4$-isogeny is already defined over $\mathbb{Q}$, we can use similar reasoning to conclude $\rho_{E^{\prime\prime},16}(G_{\mathbb{Q}})=\rho_{E^{\prime\prime},16}(G_{K})$. If it isn’t defined over $\mathbb{Q}$, [17, Proposition 3.2.] implies that the $2$-adic image $\rho_{E^{\prime\prime},2^{\infty}}(G_{\mathbb{Q}})$ equals $H_{20}$, where $H_{20}$ is the corresponding group from the list in [18]. The group $\rho_{E^{\prime\prime},2^{\infty}}(G_{K})$ is its index $3$ subgroup (it cannot be equal to it, as there would be no $2$-isogenies over $K$).
@@
However, since $H_{20}$ is defined modulo $4$, this implies that $\rho_{E^{\prime\prime},4}(G_{K})$ is an index $3$ subgroup of the reduction modulo $4$ of $H_{20}$, and $\rho_{E^{\prime\prime},2^{\infty}}(G_{K})$ is also defined modulo $4$. But then $\rho_{E^{\prime\prime},16}(G_{K})$ contains all matrices of the form $I\pmod{4}$, which is a contradiction with the starting assumption.
We conclude that $\rho_{E^{\prime\prime},16}(G_{K})=\rho_{E^{\prime\prime},16}(G_{\mathbb{Q}})$ must hold for each possible value of $s$.
Note that $\rho_{E^{\prime\prime},2^{\infty}}(G_{\mathbb{Q}})$ cannot be defined modulo $8$, as it does not contain all matrices which are congruent to $I\pmod{8}$.
It remains to search through the list of all possible $2$-adic Galois images of rational elliptic curves which are not defined modulo $8$ and see whether any of them is conjugate to a subgroup of the group of all matrices of the form
$\begin{bmatrix}x&2^{4-s}u\\ 2^{s}v&\pm 1\end{bmatrix},$
for some $s$ satisfying $0\leq s\leq 4$.
For each value of $s$, it turns out that there are no possible $2$-adic images of the desired form.
$\Box$
@@ 3.4. Eliminating $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/10\mathbb{Z}$ and $\mathbb{Z}/20\mathbb{Z}$
The groups $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/10\mathbb{Z}$ and $\mathbb{Z}/20\mathbb{Z}$ do not occur as torsion subgroups of $\mathbb{Q}$-curves over cubic number fields.
@@
Proof.   Suppose that $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/10\mathbb{Z}$ does occur for some $\mathbb{Q}$-curve $E$ over a number field $K$. We first use Lemma 2.7 with $M=2$, $N=5$ to conclude that there is a $2$-isogenous curve $E_{1}$ such that $\rho_{E_{1},20}(G_{K})$ consists (in a suitable basis) of matrices of the form $\begin{bmatrix}h&*\\ 0&*\end{bmatrix}$, where $h\in\{1,11\}$. The curve $E_{1}$ is also $K$-isogenous to an elliptic curve $E^{\prime}/K$ with a rational $j$-invariant. As before, we may assume that this isogeny is cyclic.
Now, using Corollary 2.6 with $H=\{1,11\}$, we obtain that there is an $\alpha$ dividing $20$ such that $\rho_{E^{\prime},20}(G_{K})$ is conjugate to a subgroup of $\textnormal{GL}_{2}(\mathbb{Z}/20\mathbb{Z})$ of the form
$\begin{bmatrix}h&\frac{20}{\alpha}u\\ \alpha v&z\end{bmatrix},$
with $h\in H$.
Twisting to an elliptic curve $E^{\prime\prime}$ defined over $\mathbb{Q}$, we see that $\rho_{E^{\prime\prime},20}(G_{K})$ is conjugate to a subgroup of $\textnormal{GL}_{2}(\mathbb{Z}/20\mathbb{Z})$ of the form
$\begin{bmatrix}h&\frac{20}{\alpha}u\\ \alpha v&z\end{bmatrix},$
with $h\in\{1,9,11,19\}\leq\mathbb{Z}/20\mathbb{Z}$.
If $\alpha\in\{1,4,5,20\}$, then $E^{\prime\prime}$ has a $20$-isogeny over $K$, which is not possible. Thus, $\alpha\in\{2,10\}$.
Note that in both cases $E^{\prime\prime}$ has a $5$-isogeny over $K$, so it has a $5$-isogeny over $\mathbb{Q}$ by Proposition 2.8.
Case $\alpha=2$. Denote by $J$ the group of all matrices of the form
$\begin{bmatrix}h&10u\\ 2v&z\end{bmatrix},$
with $h\in\{1,9,11,19\}\leq\mathbb{Z}/20\mathbb{Z}$.
Let $A=\rho_{E^{\prime\prime},20}(G_{K})$ and $B=\rho_{E^{\prime\prime},20}(G_{\mathbb{Q}})$. Without loss of generality, we may assume $-I\in A$.
We will use the database of all possible $2$-adic images, due to Rouse and Zureick Brown ([18]).
We do a brute force search to find all subgroups $B$ of $\textnormal{GL}_{2}(\mathbb{Z}/20\mathbb{Z})$ for which there exists a subgroup $A$ of $\textnormal{GL}_{2}(\mathbb{Z}/20\mathbb{Z})$ such that:
  • $A$ is conjugate to a subgroup of $J$,
  • $A$ is conjugate to a subgroup of $B$ of index $3$,
  • $B$ mod $5$ is conjugate to a subgroup of upper-triangular matrices,
  • $\det(A)=\det(B)=(\mathbb{Z}/20\mathbb{Z})^{\times}$,
  • $-I$ is contained in both $A$ and $B$,
  • $B$ mod $4$ is an admissible image of $\rho_{E,4}(G_{\mathbb{Q}})$ for some elliptic curve $E/\mathbb{Q}$, i.e. it belongs to the list of mod $4$ images from [18].
It turns out that there are twelve such subgroups $B$, up to conjugation. Each of the twelve subgroups is conjugate to a subgroup of one of the twelve, say $B_{12}$.
Looking at $B_{12}$ mod $4$, we see that the $2$-adic representation must be defined modulo $2$ and the mod $2$ representation is of order $3$. Such elliptic curves are parametrized by $j=x^{2}+1728$, i.e. their $j$-invariant is of the form $x^{2}+1728$ for some $x\in\mathbb{Q}$. This is the curve $X_{2}$ from the list in [18].
By Table 3 in [14], elliptic curves with a rational $5$-isogeny are parametrized by
$j=\frac{(t^{2}+10t+5)^{3}}{t},$
where $t\in\mathbb{Q}$.
Now consider the curve
$C:x^{2}+1728=\frac{(t^{2}+10t+5)^{3}}{t}.$
This is a genus $1$ curve. We can check that it has no rational points. This means that $B_{12}$ and its subgroups do not occur as images $\rho_{E,20}(G_{\mathbb{Q}})$ of rational elliptic curves, which completes the proof in the case when $\alpha=2$.
Case $\alpha=10$. The proof in this case is completely analogous to the case $\alpha=2$, as all the properties of matrices used in the brute force search are invariant under taking transposes. In the end, we end up with the same list of twelve subgroups but transposed. They are all contained in $B_{12}^{t}$. The group $B_{12}^{t}$ mod $4$ is conjugate to the same mod $4$ image from the list in [18], so we can use the exact same argument. $\Box$
@@
Proof.   Suppose that the group $\mathbb{Z}/20\mathbb{Z}$ does occur as a torsion subgroup of a $\mathbb{Q}$-curve $E/K$, where $K$ is a cubic field.
Then the image $\rho_{E,20}$ is certainly contained in the set of all matrices of the form $\begin{bmatrix}h&*\\ 0&*\end{bmatrix}$ for $h\in\{1,11\}$, but from the proof of Proposition 3.5 it follows that this is impossible.
$\Box$

### v2

@@ 3. Torsion over cubic number fields
$\mathbb{Z}/N_{1}\mathbb{Z}$ $\text{ with }N_{1}=1,\ldots,16,18,20,21,$
$\mathbb{Z}/2\mathbb{Z}\oplus\mathbb{Z}/2N_{2}\mathbb{Z}$ $\text{ with }N_{2}=1,\ldots,7.$
@@
To prove Theorem˜1.9, it remains to prove that none of the following six groups appears as torsion of a $\mathbb{Q}$-curve over a cubic field.
$\mathbb{Z}/N_{1}\mathbb{Z}$ $\text{ with }N_{1}=11,15,16,20$
$\mathbb{Z}/2\mathbb{Z}\oplus\mathbb{Z}/2N_{2}\mathbb{Z}$ $\text{ with }N_{2}=5,6.$
@@ 3.1. Eliminating $\mathbb{Z}/11\mathbb{Z}$
$K{E^{\prime}/K}E′′/.{E^{\prime\prime}/\mathbb{Q}.}$$\scriptstyle{\phi}\scriptstyle{\sim}$
@@
There are two rational $j$-invariants for which the corresponding elliptic curve $E/\mathbb{Q}$ has an $11$-isogeny over $\mathbb{Q}$. For a fixed $j$-invariant, any two rational elliptic curves with this $j$-invariant are quadratic twists. This means that we can pick any rational elliptic curve with this $j$-invariant and check that its Galois image does not have the properties it would need to have by Corollary 2.4.
The $j$-invariants for which there is an $11$-isogeny over $\mathbb{Q}$ are $11^{2}$ and $11\cdot 131^{3}$. Examples of elliptic curves with these $j$-invariants with LMFDB labels 1089.c1 and 1089.c2.
For $E^{\prime\prime}$ being any of these curves, the image of the mod 11 Galois representation $\rho_{E^{\prime\prime},11}(G_{\mathbb{Q}})$ has $220$ elements. Since $3\nmid 220$ and $\rho_{E^{\prime\prime},11}(G_{\mathbb{Q}}):\rho_{E^{\prime\prime},11}(G_{K})]$ divides $3$, we must have $\rho_{E^{\prime\prime},11}(G_{K})=\rho_{E^{\prime\prime},11}(G_{\mathbb{Q}})$.
Directly inspecting the images $\rho_{E^{\prime\prime},11}(G_{\mathbb{Q}})$, there are matrices which have neither $1$ nor 1-1 as an eigenvalue, so we reach a contradiction with Corollary 2.4. $\Box$
@@ 3.2. Eliminating $\mathbb{Z}/15\mathbb{Z}$
$K{E^{\prime}/K}E′′/.{E^{\prime\prime}/\mathbb{Q}.}$$\scriptstyle{\phi}\scriptstyle{\sim}$
@@
We will look at curves from the isogeny class 1600.i on LMFDB. Note that I-I is contained in the mod $15$ image.
By Lemma˜2.1, we have $\rho_{E^{\prime},15}\sim\psi\cdot\rho_{E^{\prime\prime},15}$, where $\psi$ is some quadratic character. Furthermore, any element of $\rho_{E^{\prime},15}(G_{K})$ has $1$ as an eigenvalue when reduced modulo $3$ and modulo $5$. This means that any element of $\rho_{E^{\prime\prime},15}(G_{K})$ either has $1$ as an eigenvalue modulo $3$ and modulo $5$, or it has 1-1 as an eigenvalue modulo $3$ and modulo $5$. Now, for each of the four curves $E^{\prime\prime}$ from the class 1600.i, we do the following:
@@
  • In each of the subgroups $H$ from the list, find a matrix $A$ such that at least one matrix among $A\pmod{3}$, $A\pmod{5}$ doesn’t have $1$ as an eigenvalue, and at least one matrix among $A\pmod{3}$, $A\pmod{5}$ doesn’t have 1-1 as an eigenvalue.
@@
      The matrices $\begin{bmatrix}12&1\\ 10&6\end{bmatrix}$ and $\begin{bmatrix}6&5\\ 5&6\end{bmatrix}$ are in $\rho_{E^{\prime\prime},15}(G_{K})$. They don’t have 1-1 as eigenvalues mod $5$, so they themselves have to be in $\rho_{E^{\prime},15}(G_{K})$. However, their product equals I-I mod $3$, which doesn’t have $1$ as an eigenvalue mod $3$, so we reach a contradiction.
@@
      Similarly as in the previous case, we end up with I-I in $\rho_{E^{\prime},15}(G_{K})$, which is a contradiction.
@@ 3.3. Eliminating $\mathbb{Z}/16\mathbb{Z}$
Proof.   Suppose that some $\mathbb{Q}$-curve $E$ over a cubic field $K$ has a point $P$ of order $16$. Then $j(E)$ is not rational.
We know that $E$ is $K$-isogenous to an elliptic curve $E^{\prime}$ with rational $j$-invariant. Since composing by an isogeny of odd degree does not effect $16$-torsion, we may assume that the isogeny $\phi:E\to E^{\prime}$ has degree $2^{k}$ for some positive integer $k$. Then $k\leq 4$ by [16, Theorem 1.2]. Suppose that $k>2$. Consider a quadratic twist $E^{\prime\prime}$ of $E^{\prime}$ such that $E^{\prime\prime}$ is defined over $\mathbb{Q}$. Then $E^{\prime\prime}$ is $2^{k}$-isogenous to a quadratic twist $E^{\prime\prime\prime}$ of $E$. However, by [16, Proposition 3.2], it follows that for $k>2$, any $2^{k}$-isogeny of a rational elliptic curve which is defined over a cubic extension is in fact defined over $\mathbb{Q}$. In particular, $E^{\prime\prime\prime}$ is also defined over $\mathbb{Q}$, so $j(E^{\prime\prime\prime})=j(E)$ is rational, a contradiction.
Thus, $k\leq 2$. Let $P$ be a $K$-rational point of order $16$ on $E$. We claim that the $K$-rational isogeny $\psi:E\to E_{2}$ with kernel $\langle P\rangle$ factors through the $2^{k}$-isogeny $\phi:E\to E^{\prime}$. Suppose for the sake of contradiction that it does not. If $k=1$, then the dual isogeny $\hat{\phi}:E^{\prime}\to E$ composed with $\psi:E\to E_{2}$ yields a cyclic $32$-isogeny of $E^{\prime}$, a contradiction with [16, Theorem 1.2]. If $k=2$ and the isogenies $\phi$ and $\psi$ are independent, then $E^{\prime}$ has a $4\cdot 16$-isogeny, a contradiction. If $k=2$ and $\phi$ and $\psi$ are not independent, then $\ker\phi\cap\ker\psi$ has size $2$. We have the following diagram, where each of the edges denotes a $2$-isogeny.
${E_{0}}E{{E^{\prime}}}{\bullet}{\bullet}$$E_{2}$}}
Since $E^{\prime}$ and $E_{2}$ are $16$-isogenous, it follows using the same argument as before that $j(E_{2})$ is rational. However, then all the elliptic curves on the path have rational $j$-invariant. Consider a rational quadratic twist $E_{0}^{\prime}$ of $E_{0}$. Then $E_{0}^{\prime}$ has two rational $2$-isogenies, so the third of its $2$-isogenies must also be rational. This implies that some twist of $E$ is a rational elliptic curve and hence $E$ has rational $j$-invariant, a contradiction. Thus, $\psi:E\to E_{2}$ factors through $\phi:E\to E^{\prime}$.
If $k=1$, then there is an $8$-isogeny $E^{\prime}\to E_{2}$, so $E_{2}$ has rational $j$-invariant. However, then $E_{2}\to E$ is a $16$-isogeny, so $E$ has rational $j$-invariant, a contradiction. Thus, $k=2$ and we have the following diagram.
${\bullet}E{{E^{\prime}}}{\bullet}$$E_{2}$}}
@@
By Corollary˜2.6, it follows that $\rho_{E^{\prime},16}(G_{K})$ is conjugate to a subgroup of the group of all matrices of the form $\begin{bmatrix}1&4x\\ 4y&z\end{bmatrix}$. Let $E^{\prime\prime}$ be a quadratic twist of $E^{\prime}$ which is defined over $\mathbb{Q}$. Then $\rho_{E^{\prime},16}(G_{K})$ is conjugate to a subgroup of the group of all matrices of the form $\begin{bmatrix}\pm 1&4x\\ 4y&z\end{bmatrix}$.
@@
We see that $E^{\prime\prime}$ has a $4$-isogeny over $K$ which is not defined over $\mathbb{Q}$. By [16, Proposition 3.2], it follows that the image of the $2$-adic Galois representation $\rho_{E^{\prime\prime},2^{\infty}}(G_{\mathbb{Q}})$ is conjugate to $H_{20}$, where $H_{20}$ is the corresponding group from the list in [17]. The group $\rho_{E^{\prime\prime},2^{\infty}}(G_{K})$ is its index $3$ subgroup (it cannot be equal to it, as there would be no $2$-isogenies over $K$).
@@
However, $H_{20}$ is defined modulo $4$. This implies that $\rho_{E^{\prime\prime},4}(G_{K})$ is conjugate to an index $3$ subgroup of the reduction modulo $4$ of $H_{20}$, and hence $\rho_{E^{\prime\prime},2^{\infty}}(G_{K})$ is also defined modulo $4$. But then $\rho_{E^{\prime\prime},16}(G_{K})$ contains all matrices of the form $I\pmod{4}$, which is a contradiction since the matrix $\begin{bmatrix}5&0\\ 0&5\end{bmatrix}$ is not conjugate to a matrix of the form $\begin{bmatrix}\pm 1&4x\\ 4y&z\end{bmatrix}$. $\Box$
@@ 3.4. Eliminating $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/10\mathbb{Z}$ and $\mathbb{Z}/20\mathbb{Z}$
The groups $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/10\mathbb{Z}$ does not occur as a torsion subgroup of $\mathbb{Q}$-curves over cubic number fields.
@@
Proof.   Suppose that $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/10\mathbb{Z}$ does occur for some $\mathbb{Q}$-curve $E$ over a number field $K$. We first use Lemma˜2.7 with $M=2$, $N=5$ to conclude that there is a $2$-isogenous curve $E_{1}$ such that $\rho_{E_{1},20}(G_{K})$ is conjugate to a subgroup of the group of all matrices of the form $\begin{bmatrix}h&*\\ 0&*\end{bmatrix}$, where $h\in\{1,11\}$. We see that $E_{1}$ has a $20$-isogeny $\phi:E_{1}\to E_{2}$ and hence does not have rational $j$-invariant by [16, Theorem 1.2]. The curve $E_{1}$ is also $K$-isogenous to an elliptic curve $E^{\prime}/K$ with a rational $j$-invariant. Let $\psi:E_{1}\to E^{\prime}$ denote this isogeny. As before, we may assume $\psi$ is cyclic of degree $2^{a}\cdot 5^{b}$ for some $a,b\geq 0$.
Since $E^{\prime}$ has a $5$-isogeny over $K$, it does not have a $4$-isogeny. Hence the degree of $\psi:E_{1}\to E^{\prime}$ is divisible by $2$ and is not divisible by $4$, so it equals $2$ or $10$. It cannot be $50$ by [16, Theorem 1.2].
Suppose that the degree is $10$. We have the following diagram, with the labels above edges being the degrees.
$E_{1}$$E_{3}$$E^{\prime}$$\scriptstyle{2}$$\scriptstyle{5}$
We claim that $E_{3}$ also has rational $j$-invariant. Consider a quadratic twist $E^{\prime\prime}$ of $E^{\prime}$ which is defined over $\mathbb{Q}$. By Proposition˜2.8, $E^{\prime\prime}$ has a rational $5$-isogeny. If this $5$-isogeny is not to a twist of $E_{3}$, then this $5$-isogenous rational elliptic curve has a $50$-isogeny over $K$ to a twist of $E_{1}$, a contradiction. Thus, $E_{3}$ is a twist of a rational elliptic curve and hence has rational $j$-invariant.
Hence, we may assume that the degree of $\psi$ is $2$. If $\psi$ does not factor through $\phi$, then $\psi$ and $\phi$ are independent and we would obtain a $40$-isogeny of $E^{\prime}$, a contradiction. Thus, $\phi$ factors through $\psi$. However, then $E^{\prime}$ is isomorphic to the starting curve $E$, since the $20$-isogeny $E_{1}\to E_{2}$ factors through both $E_{1}\to E$ and $E_{1}\to E^{\prime}$, and they are both of degree $2$. But then $E$ has rational $j$-invariant, a contradiction. $\Box$
@@
Proof.   Suppose that the group $\mathbb{Z}/20\mathbb{Z}$ does occur as a torsion subgroup of a $\mathbb{Q}$-curve $E/K$, where $K$ is a cubic field. Then there exists a $2$-isogenous elliptic curve over $K$ with torsion $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/10\mathbb{Z}$, and we already proved that this is impossible. $\Box$